OpenStudy (oswinoswald):

How do I write this in slope-intercept form? -3x - 10y = 7

OpenStudy (oswinoswald):

Please walk me through it, don't just tell me.

OpenStudy (anonymous):

just solve the equation for y

OpenStudy (teddyiswatshecallsme):

yay something i am good at, hold on and let me solve it.

OpenStudy (anonymous):

So... You start off with 4x-10y=11 Firstly, move the 4x to the other side by subtracting it. (4x-4x - 10y = 11-4x) Now, since the 4x has been moved, you are left with -10y = 11-4x. To isolate y, you have to get rid of the -10 in front of it, so divide -10 by -10 to get rid of it (And since this is Algebra, what you do on one side is also applied on the other.) Here: -10 / -10y = 11/-10 -4/-10x y = -11/10 -2/5x (Simplify the fractions from above) And make sure to shift the -2/5x to the front to match the y=mx + b format So the final equation is : y = -2/5x - 11/10 -2/5 = m (slope) x = x (:P) -11/10 = b (y-intercept) So it fits the y=mx+b equation! :) I hope this helps! :)

OpenStudy (oswinoswald):

Whoa, thanks!

OpenStudy (anonymous):

oops wrong person

OpenStudy (oswinoswald):

Still helped haha

OpenStudy (anonymous):

who cares u get the concept

OpenStudy (anonymous):

lol

OpenStudy (teddyiswatshecallsme):

So to start off, slope intercept is in form y = mx+b right? And this is in standard form, what wee need to do is get 3x over t o the other side of the equal sign, since it is a -3x you do reverse operations or inverse operations, and add 3x to both sides of the equation, according to properities of equality, once you do that you get -10y = 3x + 7 and now you need to isolate or solve for y, how you do this is you use inverse operations with the 10y... 10y is basically saying 10 times y so, the opposite of mulitplication is division so you take the -10y and you divide each expression in the equation by -10, so the equation then becomes, y = -3x/5 + -7/10, 10 simplifies down to 5 if you are wondering where i got 5 from..... I may be incorrect but that is what i believe is the answer...

OpenStudy (anonymous):

...................